OpenStudy (anonymous):

Solve the inequality for y. ≤−67y10 Simplify your answer as much as possible.

OpenStudy (anonymous):

@armmonipowtoon helps please

OpenStudy (anonymous):

@TheSmartOne how do I do this question as well

TheSmartOne (thesmartone):

Could you write your question using the equation editor or the draw tool? .-.

TheSmartOne (thesmartone):

inequality**** not equation .-.

OpenStudy (anonymous):



OpenStudy (anonymous):

sorry if it's sloppy kind of hard to draw on this

OpenStudy (anonymous):

@TheSmartOne

TheSmartOne (thesmartone):

To solve it, first multiply by the reciprocal. What is the reciprocal of \(\sf\Large\frac{-6}{7}\) Hint: Reciprocal of \(\sf\Large\frac{a}{b}\) is \(\sf\Large\frac{b}{a}\)

TheSmartOne (thesmartone):

and note that since you will be multiplying by a negative number on both sides, the sign will flip. So the \(\sf \le\) will become a \(\sf \ge\)

OpenStudy (anonymous):

oh it would be 7 over -6

TheSmartOne (thesmartone):

correct, and so multiply both sides by that and don't forget the switch the sign :)

OpenStudy (anonymous):

so y flip the sign 70 over -6?

TheSmartOne (thesmartone):

correct \(\sf\large y \ge -\frac{70}{6}\) One last step, just simplify -70/6 so it's in simplest form